# FastGS: Training 3D Gaussian Splatting in 100 Seconds

### v2

@@ 3 Background: 3D Gaussian Splatting
Figure 3: The pipeline of FastGS. (a) We redesign the ADC of the vanilla 3DGS [kerbl3Dgaussians] based on multi-view consistency. To accurately assess the importance of each Gaussian, we sample training views and generate the corresponding per-pixel L1 loss maps. For each sampled view, a multi-view score is computed for each Gaussian by counting the number of high-error pixels within its 2D footprint, which is subsequently used to guide Gaussian densification and pruning. (b) Taming-3DGS [mallick2024taming] primarily computes the importance score based on Gaussian-associated properties across sampled views. (c) Speedy-Splat [hanson2025speedy] computes the Gaussian score by accumulating Gaussian-associated Hessian approximations across all training views. We visualize the densification results from 0.5K to 15K iterations without any pruning on the far left of the figure, and the pruning results after removing Gaussians at the 6K-th iteration on the far right.
@@ 4.1 Overview
The framework of our method is shown in Fig. 3. We initialize 3D Gaussians using SfM point clouds and train the 3DGS model on multi-view images. The addition and removal of 3D Gaussians are controlled via the proposed multi-view consistent densification and pruning strategies. As illustrated in Fig. 3 and Fig. 3, Taming-3DGS [mallick2024taming] and Speedy-Splat [hanson2025speedy] also estimate importance scores from multiple views, for densification and pruning, respectively. However, both rely on Gaussian-associated scores to control the number of Gaussians. This suboptimal use of multi-view information results in redundancy for Taming-3DGS [mallick2024taming] and degraded rendering quality for Speedy-Splat [hanson2025speedy]. In contrast, as illustrated in Fig. 3, our method evaluates the importance of each Gaussian based on multi-view reconstruction quality, rather than gaussian-associated properties. Furthermore, our method leverages multi-view consistency constraints to effectively guide both densification and pruning, which will be detailed in Sections Sec. 4.2 and Sec. 4.3. To further improve rasterization efficiency, Sec. 4.4 introduces the compact box (CB) we use.
@@ 4.4 Compact Box
During the preprocessing stage of rasterization, the vanilla 3DGS [kerbl3Dgaussians] uses the 3-sigma rule to obtain 2D ellipses, generating many Gaussian–tile pairs that introduce computational redundancy and reduce rendering efficiency. Speedy-Splat [hanson2025speedy] partially addresses this with precise tile intersection, yet we observe that some 2D Gaussians still have a negligible impact on pixels in certain tiles. As illustrated in Fig. 4, to further reduce unnecessary pairs, we propose compact box, which prunes Gaussian–tile pairs with minimal contribution based on Mahalanobis distance from the Gaussian center, thereby accelerating rendering while maintaining quality. Details are provided in Sec. 8 of the supplementary material.
@@ 8 Details of Compact Box
During the preprocessing stage of rasterization, the vanilla 3DGS [kerbl3Dgaussians] employs the 3-sigma rule to coarsely obtain effective 2D ellipses, which results in a large number of Gaussian–tile pairs as shown in Fig. 7, introducing computational redundancy and significantly reducing rendering efficiency. To address this issue, Speedy-Splat [hanson2025speedy] proposes the precise tile intersect method to reduce the number of Gaussian–tile pairs. However, according to our observations, there remains room for further improvement, as the 2D Gaussians have a negligible impact on the pixels within the numbered tiles in Fig. 7. Therefore, we propose a method named Compact Box that further reduces the 2D Gaussian content to eliminate unnecessary Gaussian–tile pairs, as shown in Fig. 7.

### v3

@@ 3 Background: 3D Gaussian Splatting
Figure 3: The pipeline of FastGS. (a) We redesign the ADC of the vanilla 3DGS [kerbl3Dgaussians] based on multi-view consistency. To accurately assess the importance of each Gaussian, we sample training views and generate the corresponding per-pixel L1 loss maps. For each sampled view, a multi-view score is computed for each Gaussian by counting the number of high-error pixels within its 2D footprint, which is subsequently used to guide Gaussian densification and pruning. (b) Taming-3DGS [mallick2024taming] primarily computes the importance score based on Gaussian-associated properties across sampled views. (c) Speedy-Splat [hanson2025speedy] computes the Gaussian score by accumulating Gaussian-associated Hessian approximations across all training views. We visualize the densification results from 0.5K to 15K iterations without pruning on the far left, and pruning results on the far right using Speedy-Splat [hanson2025speedy]’s pruning strategy and VCP on vanilla 3DGS [kerbl3Dgaussians].
@@ 4.1 Overview
The framework of our method is shown in Fig. 3. We initialize 3D Gaussians using SfM point clouds and train the 3DGS model on multi-view images. The addition and removal of 3D Gaussians are controlled via the proposed multi-view consistent densification and pruning strategies. As illustrated in Fig. 3 and Fig. 3, Taming-3DGS [mallick2024taming] and Speedy-Splat [hanson2025speedy] also estimate importance scores from multiple views, for densification and pruning, respectively. However, both rely on Gaussian-associated scores to control the number of Gaussians. This suboptimal use of multi-view information results in redundancy for Taming-3DGS [mallick2024taming] and degraded rendering quality for Speedy-Splat [hanson2025speedy]. In contrast, as illustrated in Fig. 3, our method evaluates the importance of each Gaussian based on multi-view reconstruction quality, rather than gaussian-associated properties. Furthermore, our method leverages multi-view consistency constraints to effectively guide both densification and pruning, which will be detailed in Sections Sec. 4.2 and Sec. 4.3. To further improve rasterization efficiency, Sec. 4.4 introduces the compact box (CB) we use, which is adapted from the precise tile-intersection strategy proposed in Speedy-Splat [hanson2025speedy].
@@ 4.4 Compact Box
During the preprocessing stage of rasterization, the vanilla 3DGS [kerbl3Dgaussians] uses the 3-sigma rule to obtain 2D ellipses, generating many Gaussian–tile pairs that introduce computational redundancy and reduce rendering efficiency. Speedy-Splat [hanson2025speedy] partially addresses this with precise tile intersection, yet we observe that some 2D Gaussians still have a negligible impact on pixels in certain tiles. As illustrated in Fig. 4, to further reduce unnecessary pairs, we introduce a compact box (CB), which builds upon and extends Speedy-Splat [hanson2025speedy]’s precise tile-intersection strategy by pruning Gaussian–tile pairs with minimal contribution based on the Mahalanobis distance from the Gaussian center. This further accelerates rendering while maintaining quality. Details are provided in Sec. 8 of the supplementary material.
@@ 8 Details of Compact Box
During the preprocessing stage of rasterization, vanilla 3DGS [kerbl3Dgaussians] employs the 3-sigma rule to coarsely obtain effective 2D ellipses, which results in a large number of Gaussian–tile pairs as shown in Fig. 7, introducing computational redundancy and significantly reducing rendering efficiency. To address this issue, Speedy-Splat [hanson2025speedy] proposes a precise tile-intersection method to reduce the number of Gaussian–tile pairs. However, according to our observations, there remains room for further improvement, as some 2D Gaussians still have a negligible impact on the pixels within the numbered tiles in Fig. 7. Therefore, we propose Compact Box, which builds upon and refines Speedy-Splat’s precise tile-intersection strategy to further reduce the effective 2D Gaussian region and eliminate unnecessary Gaussian–tile pairs, as illustrated in Fig. 7.
@@
In implementation,  Eq. 15 is integrated into Speedy-Splat [hanson2025speedy]’s snugbox, where the parameter $\beta$ further reduces the 2D Gaussian footprint and shrinks its intersection region with tiles, thus forming our CB. We then obtain the Gaussian–tile pairs using Speedy-Splat [hanson2025speedy]’s accutile method. The comparison between snugbox [hanson2025speedy] and our CB can be found in Tab. 16. We sincerely appreciate the excellent work of Speedy-Splat [hanson2025speedy].
@@ 11 Additional Ablation
Table 16: Comparison of rasterization methods. Experiments are performed on the Mip-NeRF 360 dataset [barron2022mip] with 3DGS-accel [kerbl3Dgaussians, mallick2024taming] as the baseline.
[…]
3DGS-accel 7.10 27.46 0.810 0.226 2.64M 182
+snugbox [hanson2025speedy] 6.50 27.46 0.811 0.224 2.69M 288
+CB. 6.13 27.44 0.810 0.223 2.78M 303